OpenStudy (anonymous):

On July 4th, dizzyland had 67,000 visitors and collected approximately $2,814,000. how much money should dizzyland expect to receive on New Years Day, when park attendance reaches 71,000 people?

OpenStudy (noseboy908):

Are you familiar with rates, like "miles PER hour"?

OpenStudy (anonymous):

yes

OpenStudy (noseboy908):

alright, then you know that rates are simply (something) PER (something else). In this case, it's dollars per person, as you need to solve for how much each person paid for their ticket. So, how would you solve for the cost per ticket?

OpenStudy (anonymous):

A proportion?

OpenStudy (noseboy908):

Yes. So, any ideas about my last question's answer?

OpenStudy (anonymous):

y = kx 2814000 = k(67000) k = 42 y = kx y = 42*71000 y = 2982000 $2,982,000 is the answer?

OpenStudy (noseboy908):

Yes!

OpenStudy (anonymous):

Thank you (:

OpenStudy (noseboy908):

You are most welcome! If I was helpful, please hit that "best response" button. I really appreciate it!

OpenStudy (noseboy908):

thanks! Any more problems you need help with?

OpenStudy (anonymous):

No I'm good. Thanks for asking though!

OpenStudy (noseboy908):

Feel free to mention me in a reply like this one in the future if you need my help. It looks like "@noseboy908", if you don't know. With that, I directly receive an email that you mentioned me. (Actually, could you test that out for me? Just type my name as I showed you.)

OpenStudy (anonymous):

@noseboy908

OpenStudy (noseboy908):

Hmm, for some reason, that didn't send me anything. I'll ask a mod about it in a sec. Keep this question up for a minute more please.

OpenStudy (noseboy908):

Now, once more, please mention me again, I think I got it so it will notify me.

OpenStudy (anonymous):

@noseboy908

OpenStudy (noseboy908):

Alright, it's set. It will email me whenever you, or anyone else, mentions me if i'm offline. So, if you ever need help, just save my name to mention me, and I'll help ASAP.

OpenStudy (anonymous):

Cool. Thanks again!

OpenStudy (noseboy908):

Welcome, I'm glad to be of service.